OpenStudy (anonymous):

ALGEBRA 1 HELP EASYYYY please help and explain 3. The length of a rectangle is 7 mm longer than its width. Its perimeter is more than 62 mm. Let w equal the width of the rectangle. a. Write an expression for the length in terms of the width. b. Use these expressions to write an inequality based on the given information. c. Solve the inequality, clearly indicating the width of the rectangle.

OpenStudy (michele_laino):

From your data we can write: \[l=7+w\] where l is length, and w is width

OpenStudy (anonymous):

ok so that would be the expression?

OpenStudy (michele_laino):

please, try to express the perimeter in function of l and w

OpenStudy (anonymous):

How would I do b?

OpenStudy (michele_laino):

\[2p=2*\left( l+w \right)\] where 2p is the perimeter

OpenStudy (anonymous):

ok but where would the 62 come into play?

OpenStudy (michele_laino):

from the text of your problem, we can write: \[2p >62\] now, please, try to substitute the formula above for 2p

OpenStudy (anonymous):

Im so confused! Can anyone else help!!

OpenStudy (michele_laino):

\[2\times \left( l+w \right)>62\]

OpenStudy (anonymous):

@StarJ9 can you help?

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

I don't see how anyone else would explain this better then what Michele posted. Remember that the perimeter of a rectangle is p= w+w + l+l, or p = 2w + 2l, there are two widths, and two lengths. So, from above 2(w+l) he just factor the two out. The perimeter is more then 62, so p > 62. We know what the perimeter is 2w+2l. So, 2w + 2l > 62. The length of a rectangle is 7 mm longer than its width. So, the length is the width + 7 more. length = 7 + w Putting it all together is 2w + 2(7+w) > 62

OpenStudy (anonymous):

Solving for w gives us 12, and plugging w into l = 7 + (12) = 19. w = 12 l = 19 2(12) + 2(19) = 62 ...something is wrong...62 is not > 62...